OpenStudy (anonymous):

Solve for d. -10 = -2 + 1/2 d

OpenStudy (anonymous):

+2 on both side = -8=1/2d, then divide by 1/2 both side to get d

OpenStudy (anonymous):

-16=d

OpenStudy (anonymous):

-10 = -2 + 1/2 d First +2 on both side, -10+2 =-8 and -2+2=0 so the resulting equation is: -8=1/2d Then to get d by itself you need to divide by 1/2. So -8/0.5 = -16 and .5/.5=0 so answer is -16=d :D